OpenStudy (idealist10):

Solve for y: y^2/2-2y=x^3+x^2+x+C

OpenStudy (idealist10):

I multiplied the whole equation by 2, I got y^2-4y=2(x^3+x^2+x+C)

OpenStudy (anonymous):

\[\frac{y^2}{2}-2y=\cdots~~?\] Try completing the square.

OpenStudy (idealist10):

Sorry, I'm not good at that.

OpenStudy (anonymous):

Okay, well we don't really have to worry about the right hand side, so I'll just focus on the left and you can take it from there. \[\frac{1}{2}y^2-2y=\frac{1}{2}\left(y^2-4y\right)\] Add and subtract 4: \[\frac{1}{2}\left(y^2-4y+4-4\right)\] Factor the first three terms: \[\frac{1}{2}\left((y-2)^2-4\right)\] Distribute the 1/2: \[\frac{1}{2}(y-2)^2-2=x^3+\cdots\]

OpenStudy (anonymous):

Now you can isolate the \(y\) to get an explicit solution. Don't forget the \(\pm\sqrt{\cdots}\).

OpenStudy (idealist10):

HAHA, didn't learn the completing square method fully in algebra before. Sorry. :)

OpenStudy (idealist10):

Got it! But I don't know why the answers in the book are slightly different. I guess they're all wrong?

OpenStudy (anonymous):

Two possibilities that I can see. (1) When you add 2 to the right side, then multiply by 2, it could be that the 2 you add is absorbed into \(C\). Or, (2) it's not absorbed into \(C\)... I think (1) is more likely. What does your book say?

OpenStudy (idealist10):

My book says y=2+/-sqrt(2(x^3+x^2+x+c)) but I got y=2+/-sqrt(2(2+x^3+x^2+x+c)).

OpenStudy (idealist10):

So what's the correct answer then?

OpenStudy (anonymous):

Looks like it does get absorbed. I'd go with the book's answer.

OpenStudy (idealist10):

Okay. :)

OpenStudy (anonymous):

From Mathematica 9:\[y=2\pm \sqrt{2} \sqrt{c+x^3+x^2+x} \]